OpenStudy (anonymous):

The integral of 7((cos(x)+sin(x))/(sin(2x))

OpenStudy (anonymous):

hint: sin(2x) = 2sin(x)cos(x)

OpenStudy (anonymous):

cant you also use sin2x=1/2 cos(2x)

OpenStudy (anonymous):

how is sin(2x) = (1/2) cos(2x)?

OpenStudy (anonymous):

that trig identity isn't true

OpenStudy (anonymous):

oh i think it is (cos(2x)/2) is that not right ?

OpenStudy (anonymous):

for integration by parts?

OpenStudy (anonymous):

how about this, 7(cosx + sinx)/(2sinx cosx) = (7/2) [ cosx/(sinx cosx) + sinx/(sinx cosx) ] = (7/2) (cscx + secx) can you finish from here?

OpenStudy (anonymous):

i just gave you 95% of the answer

OpenStudy (anonymous):

okay ill see if i can finish it

OpenStudy (anonymous):

how do i figure out what the integral of the csc and sec are ?

OpenStudy (anonymous):

7/2((ln(tan(x)+sec(x))-ln(csc(x)+cot(x)))

OpenStudy (anonymous):

well, you can look up the basic integral for trig and memorize them. Just like you memorize that the integral of cos(x) is sin(x). The integral of cscx is ln(cscx - cotx)